Fermi National Laboratory


Accelerator Update

Friday March 12
The day shift began with Operations monitoring store 3289, with the Antiproton Source (Pbar) stacking, and with the Main Injector (MI), Recycler, Switchyard (SY), and MiniBooNE taking beam and conducting studies.

Operators saw some beam irregularities at 8:28 AM that required them to tune beam for the Collider Detector at Fermilab (CDF) and D-zero (D0) experiments.

At 11:22 PM, operators interlocked the Meson Test (MTest) enclosures.

Saturday March 13
The midnight shift began with Operations monitoring store 3289 and the Recycler stash, with Pbar stacking, and with beam to SY and MiniBooNE.

At 12:17 AM, a Linac RF station (LRF2) tripped off. It reset okay.

Operations terminated store 3289 at 12:51 AM. Two Tevatron (TeV) experts began an orbit study at 1:20 AM.

At 2:11 AM, while attempting to abort the study beam, the TeV suffered a sector A1 quench. Pbar resumed stacking.

Operations put the TeV through a dry squeeze at 3:59 AM.

Operations put the TeV in shot setup at 4:22 AM and at 6:28 AM established store 3291. Its initial luminosity was 62.99E30. Pbar resumed stacking.

Booster tripped off at 6:37 AM due to a bad cable going to a beam detector. (This halted all downstream beam.) Operators called in a safety system technician to make repairs.

The safety system tech finished his repair at 9 AM. Stacking, MiniBooNE beam, and beam to MTest all resumed.

Operations began sending beam to Meson Center (MCenter) at 9:30 AM.

At 2 PM, Recycler began stashing transfers of antiprotons from Pbar. They had stashed four shots by 2:20 PM. Stacking then resumed.

A Booster RF station's output went to zero at 3:19 PM. Operators investigated and had it back up about forty minutes later.

MTest experimenters finished using the beam at 9:47 PM.

A Linac RF station (LRF4) tripped off at 9:57 PM. It reset okay.

Sunday March 14
The midnight shift began with Operations monitoring store 3291, with Pbar stacking, with MI and Recycler conducting studies, and with beam to SY and MiniBooNE.

Operations terminated store 3291 at 6 AM.

At 6:16 AM, Operation put the TeV through a dry squeeze and then into shot setup at 6:41 AM.

Operations established store 3293 at 9:17 AM with an initial luminosity of 51.87E30. Pbar resumed stacking.

By 10:01 AM, a Linac RF station (KRF5) had tripped off three times. It reset okay.

At 10:23 AM, beam to Meson tripped off. Operators investigated and found a MCenter power supply (MC1D) leaking LCW and a M01 cooling system off. Operators, with the help of an expert, got this system back on, but due to the problem with MC1D experts decided that MCenter would stay off until after the shutdown. MC1D repairs will occur on Monday.

Operations put the Recycler, MI, and Pbar into shot setup for antiproton transfers at 4:20 PM. Recycler finished stashing and Pbar resumed stacking at 5:23 PM.

Two Booster RF stations (BRF7 & 8) tripped off at 6:22 PM due to high vacuum pressure. They both reset okay.

At 8:17 PM, MI, TeV, Pbar, and Booster experts began studies.

Pbar studies ended at 10:22 PM.

MTest finished using beam at 11:23 PM.

Recycler studies finished at 11:24 PM.

Operations terminated store 3293 at 11:34 PM. Operations prepared the TeV for the upcoming safety system tests.

Monday March 15
The midnight shift began with no beam in any Fermilab accelerators and Operations preparing areas for the safety system tests and the ten-day shut down.

Safety system techs began their TeV tests at 12:16 AM. They finished at 1:31 AM and moved on to other areas.

Vacuum techs accessed the TeV at sector A4 to begin repairs at 2:16 AM.

Safety system techs completed all their system tests at 3:42 AM.

At 7 AM, high voltage techs deenergized feeder 24 for thirty minutes for maintenance.

Today's Plans
Today is the beginning of a ten-day shut down for the Fermilab complex. The main reason for this shutdown is to repair the TeV A4 vacuum. This requires spoiling the vacuum in this area and warming up the cryo system. All the other machines will conduct maintenance and upgrades during the first five days of the shutdown and then startup next weekend.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/15/2004   email Fermilab